### v1

@@ 10.1 Two Two-Parameter Toy Models
Both toy models initialise two tip EC along the bottom of a the square domain $\mathcal{I}=[0,1]^{2}$, which move diagonally upwards. The tip ECs anastomose, branch, and then anastomose again, forming a central loop before reaching the VEGF source at the top of the domain. We derive two stochastic spatial models from this simple construction by parametrising the size and shape of the central loop. The first model, Toy-Circle (TC), has two parameters $r$ and $c$ which define the centre and radius of a circular loop in the middle of the domain. Given a parameter pair $(r,c)$, the TC model produces a blood vessel network whose central loop is a circle of radius $r+\varepsilon$ and centre $(0.5,c+\varepsilon)$, where each $\varepsilon$ is identically and independently drawn from $\mathcal{U}_{[-0.03,0.03]}$, the uniform distribution between $-0.03$ and $0.03$. The randomly sampled $\varepsilon$ are added to each parameter to introduce random variation into the model which is simple enough to allow derivation of the model’s likelihood function. The second model, Toy Ellipse (TE), uses a similar construction to produce a vascular network. However, the TE model uses parameters $a$ and $b$, to define the size of the horizontal radius and vertical radius of an ellipse which makes up the loop in the centre of the domain. Given a parameter pair $(a,b)$, the TE model produces a blood vessel network whose central loop is an ellipse centred at $(0.5,0.5)$ with horizontal radius $a+\varepsilon$ and vertical radius $b+\varepsilon$. The noise $\varepsilon$ is sampled in the same way as in the TC model. Each toy model, including the range of each parameter, is summarised in Table 2.
@@ 10.3 Parameter Inference and Model Selection
To test our method of spatial model selection, we perform Step 3 of Section 3 of the main text and approximate the model posterior $p(m|\mathcal{D}^{*})$ for the same four test-cases, giving results in Figure 9. Since test-cases 1 and 2 of the TC model contain circles not centred at $(0.5,0.5)$, and test-cases 1 and 2 of the TE model contain ellipses with unequal horizontal and vertical radii, only the true model can produce their observed data. In these test-cases, we successfully infer the true model and correctly approximate posterior probabilities as $1$ for the true model and 00 for the other model. The observed data in test-cases 3 and 4 of each toy model, however, could have been generated by either model. In these test-cases, the central loop in the observed data is is either a circle with centre $c=(0.5,0.5)$, or an ellipse with equal horizontal and vertical radii, which either toy model can reproduce. We therefore expect a non-zero posterior probability for each model in these test-cases, and we successfully approximate this. The approximate model posterior is still able to identify the true model in these cases, and we estimate only a small posterior probability of the incorrect mode in each test-case. While both models can simulate data similar to the observed data in each test-case, the true model will do so more often (for more parameter values). The training data from which we learn the relationship between summary statistics and model index therefore contains more simulations similar to the observed data when the true model is used, which may be why it predicts the true model with higher probability.
@@
Acknowledgements
Funding
@@ Acknowledgements
All code is available at https://github.com/rmcdomaths/tms/ Code for the AC model was taken from Nardini et al. (2021) and https://github.com/johnnardini/Angio_TDA. The ABCpy python package https://github.com/eth-cscs/abcpy was used to run the ABC-SMC algorithm. Parts of Figure 1 were created in https://BioRender.com

### v2

@@ 10.1 Two Two-Parameter Toy Models
Both toy models initialise two tip EC along the bottom of a the square domain $\mathcal{I}=[0,1]^{2}$, which move diagonally upwards. The tip ECs anastomose, branch, and then anastomose again, forming a central loop before reaching the VEGF source at the top of the domain. We derive two stochastic spatial models from this simple construction by parametrising the size and shape of the central loop. The first model, Toy-Circle (TC), has two parameters $r$ and $c$ which define the centre and radius of a circular loop in the middle of the domain. Given a parameter pair $(r,c)$, the TC model produces a blood vessel network whose central loop is a circle of radius $r+\varepsilon$ and centre $(0.5,c+\varepsilon)$, where each $\varepsilon$ is identically and independently drawn from $\mathcal{U}_{[-0.03,0.03]}$, the uniform distribution between 0.03-0.03 and $0.03$. The randomly sampled $\varepsilon$ are added to each parameter to introduce random variation into the model which is simple enough to allow derivation of the model’s likelihood function. The second model, Toy Ellipse (TE), uses a similar construction to produce a vascular network. However, the TE model uses parameters $a$ and $b$, to define the size of the horizontal radius and vertical radius of an ellipse which makes up the loop in the centre of the domain. Given a parameter pair $(a,b)$, the TE model produces a blood vessel network whose central loop is an ellipse centred at $(0.5,0.5)$ with horizontal radius $a+\varepsilon$ and vertical radius $b+\varepsilon$. The noise $\varepsilon$ is sampled in the same way as in the TC model. Each toy model, including the range of each parameter, is summarised in Table 2.
@@ 10.3 Parameter Inference and Model Selection
To test our method of spatial model selection, we perform step 3 of section 3 of the main text and approximate the model posterior $p(m|\mathcal{D}^{*})$ for the same four test-cases, giving results in Figure 11. Since test-cases 1 and 2 of the TC model contain circles not centred at $(0.5,0.5)$, and test-cases 1 and 2 of the TE model contain ellipses with unequal horizontal and vertical radii, only the true model can produce their observed data. In these test-cases, we successfully infer the true model and correctly approximate posterior probabilities as $1$ for the true model and 0 for the other model. The observed data in test-cases 3 and 4 of each toy model, however, could have been generated by either model. In these test-cases, the central loop in the observed data is is either a circle with centre $c=(0.5,0.5)$, or an ellipse with equal horizontal and vertical radii, which either toy model can reproduce. We therefore expect a non-zero posterior probability for each model in these test-cases, and we successfully approximate this. The approximate model posterior is still able to identify the true model in these cases, and we estimate only a small posterior probability of the incorrect mode in each test-case. While both models can simulate data similar to the observed data in each test-case, the true model will do so more often (for more parameter values). The training data from which we learn the relationship between summary statistics and model index therefore contains more simulations similar to the observed data when the true model is used, which may be why it predicts the true model with higher probability.
@@
11 Supplementary Results
We experiment with different values of $t_{\text{final}}$ to investigate whether our pipeline can infer parameters and select the correct model with less simulation information. Specifically, we modify the AC, SL, and PS models so that they terminate after times $t_{\text{final}}=2.5,2.5,1.5$ respectively, meaning that ECs only move around half of the way to the tumour within the simulation.
Figure 12 shows the results of parameter inference using the same test-cases as Figure 3 of the main text (but with simulation times decreased). Interestingly, the non-branching parameters in each model are inferred as well as, if not better than, when the simulations are run for the full time period. The approximate posterior for these parameters are narrower than in Figure 3 of the main text. When the simulation time is shorter, the movement parameters (chemotaxis in all models, and either haptotaxis, randomness and turning rate) appear to have a stronger effect on the simulated data. For example, the value of the chemotaxis parameter determines how quickly ECs travel towards the tumour. In simulations which run up to $t_{\text{final}}=4$, most ECs reach the tumour. However when time is limited, the chemotaxis parameter affects how close ECs get to the tumour, which may be why its value is inferred with more certainty when simulation time is limited.
The branching parameters $a_{\text{br}}$ and $c_{\text{br}}$ are inferred less accurately (and approximate parameter posterior distributions are wider) when simulations are time-limited, and are in some cases not inferred correctly. Perhaps unsurprisingly, when simulations are run for shorter times, the ECs exhibit less branching behaviour, which may explain why these parameters are now more difficult to infer.
Figure 13 attempts model selection, again using the same (modified) test-cases as Figure 4 of the main text. In each test-case, the correct model is identified, despite the simulations being time-limited. However, the posterior probability of the incorrect models is higher, and it appears to be more difficult to identify the correct model when less simulation information is available. The ‘model predictions’ appear to suffer from the shorter simulation time. Since some parameters (in particular, those which regulate branching) are now uncertain, parameters sampled from each model’s approximate posterior distribution now simulate data that are significantly visually different to the observed data.
Figure 12: As in Figure 3 of the main text, we infer each parameter in the AC, SL, and PS models, this time modifying each model (and each test-case) to simulate EC movement for a shorter time period. We again simulate each model $10$ times at known parameter values to generate two synthetic test-cases for each model, and show the final time-step of one such simulation. We project the approximate ABC-SMC posterior to each parameter pair and plot the resulting distributions (fitting a Gaussian kernel to the parameter values accepted in the final population of the ABC-SMC algorithm). We also plot the true parameter which generated the test-case.
Figure 13: As in Figure 4 of the main text, we approximate the model posterior, again modifying each model (and each test-case) to simulate EC movement for a shorter time period. We again show one ‘prediction’– an example of data simulated using an inferred parameter from each model’s approximate parameter posterior. Now that some of the approximate parameter posteriors are so uncertain, some model predictions are very different from the observed data.
Funding Acknowledgements
@@ 11 Supplementary Results
All code is available at https://github.com/rmcdomaths/tms/ and archived at https://doi.org/10.5281/zenodo.17392787. Parts of figures 1 and 2 were created in Created in BioRender https://BioRender.com/.